one angle of a triangle is twice the first angle. The third angle is eight more that the first angle. find the measure of each angle.
Answer by algebrahouse.com(1659)

"one angle of a triangle is twice the first angle. The third angle is eight more that the first angle. find the measure of each angle."

x = first angle
2x = second angle {one angle is twice the first}
x + 8 = third angle {third angle is 8 more than first}

x + 2x + x + 8 = 180 {angles of a triangle add up to 180}
4x + 8 = 180 {combined like terms}
4x = 172 {subtracted 8 from both sides}
x = 43 {divided both sides by 4}
2x = 86 {substituted 43, in for x, into 2x}
x + 8 = 51 {substituted 43, in for x, into x + 8}

43, 86, and 51 are the three angles
www.algebrahouse.com